OpenStudy (anonymous):

let (theta sign) be an angle in quadrant II such that sin(theta sign)=3/4 find the exact values of sec (theta sign) and cot (theta sign)

OpenStudy (anonymous):

\[\cos = \sqrt{(1-\sin^2x)}\] and sec = 1/cos = 4/sqrt(5)

OpenStudy (anonymous):

\[\cos = \sqrt{(1-\sin^2x)}\] and sec = 1/cos = 4/sqrt(7) cot = cos/sin = sqrt(7)/3

OpenStudy (anonymous):

Since this is quadrant two, sec and cot are negative.

OpenStudy (anonymous):

p.s.gabrielle:you had a logarithms question earlier sorry i wasn't able to help then...if you still need help i would be glad to see if i could be of assistance